A jet travels 530 miles in 5 hours. At this rate how far could the jet fly in 11 hours? What is the rate of speed of the jet?

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

530/5 = x/11
x = 11*530/5
x = 1166 miles
---
speed = 530/5
speed = 106 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by Alan3354(69443)

There is no "rate of speed."
It's either rate, or speed.
===================
Confirmed by the Dept of Redundancy Dept.